# Exploration of Masked and Causal Language Modelling for Text Generation

### v1

@@ Abstract
Large Language Models (LLMs) have revolutionised the field of Natural Language Processing (NLP) and have achieved state-of-the-art performance in practically every task in this field. However, the prevalent approach used in text generation, Causal Language Modelling (CLM), which generates text sequentially from left to right, inherently limits the freedom of the model, which does not decide when and where each token is generated. In contrast, Masked Language Modelling (MLM), primarily used for language understanding tasks, can generate tokens anywhere in the text and any order. This paper conducts an extensive comparison of MLM and CLM approaches for text generation tasks. To do so, we pre-train several language models of comparable sizes on three different datasets, namely 1) medical discharge summaries, 2) movie plot synopses, and 3) authorship verification datasets. To assess the quality of the generations, we first employ quantitative metrics and then perform a qualitative human evaluation to analyse coherence and grammatical correctness. In addition, we evaluate the usefulness of the generated texts by using them in three different downstream tasks: 1) Entity Recognition, 2) Text Classification, and 3) Authorship Verification. The results show that MLM consistently outperforms CLM in text generation across all datasets, with higher quantitative scores and better coherence in the generated text. The study also finds no strong correlation between the quality of the generated text and the performance of the models in the downstream tasks. With this study, we show that MLM for text generation has great potential for future research and provides direction for future studies in this area.
@@ I Introduction
The rise of Large Language Models (LLMs) has gained much attention in recent years. These models have demonstrated the remarkable ability of complex reasoning and generate human-like text. However, it became apparent since its very introduction that the transformer architecture, the backbone of LLMs, has numerous limitations. In fact, the attention mechanism [1], the core component of transformers, struggles with capturing long-range dependencies and does not scale efficiently to capture longer sequences. This finding led researchers to explore various ways to overcome these limitations and introduce many transformer variants [2, 3, 4]. This collective effort also led to the exponential growth in size of these models.
@@ 5 The quick brown fox jumps over the lazy dog
This process is called Causal Language Modelling (CLM) because the probability of generating each next token is conditioned only on the tokens that appear before it in the sequence. This unidirectional approach inevitably restricts their capacity to leverage the full context of a text, as these models are prevented from incorporating or generating tokens that appear later in the sequence. This limitation is not present when using another approach called Masked Language Modelling (MLM).
@@
Despite their potential, MLM applications in text generation, particularly for tasks such as synthetic data generation [7], have been greatly overlooked, leaving their generative capabilities massively unexplored. In this paper, we aim to challenge this assumption in text generation by comparing CLM with MLM using models of comparable size. By exploring this potential, we aim to find new directions for this NLP area and understand under what conditions MLM for text generation can shine. Since there is little research on this topic, our starting point will be synthetic data generation, as we found that this approach would be easier to evaluate since we can easily compare the synthetic dataset with the original one and also use the synthetic data to train models in downstream tasks and evaluate their performance.
@@
    Does domain-specific knowledge used for pre-training and fine-tuning matter for generating text that is of higher quality with regard to the quantitative metrics?
@@
    Does higher quality in the texts equal better performance in downstream tasks?
@@
    What kind of tokens provide enough context to generate high-quality text?
To answer these questions, we aimed to thoroughly evaluate and compare the ability of CLM and MLM models of comparable sizes to generate text. The following are the main tasks of this investigation:
@@
    Fine-tune a number of Pre-trained Language Models (PLMs), including T5, BART, and BERT and their variations.
@@
    Compare the generations of MLM and CLM models with standard quantitative metrics and qualitatively assess the generations.
@@
    Examine the usefulness of the generations of the fine-tuned models in downstream tasks, namely Named Entity Recognition (NER), Text Classification and Authorship Verification.
@@
    Interpret the results of the evaluations and conclude the optimal model types and masking strategies for different text generation tasks across the three different datasets.
@@ II Background and Related Work
II-A Generative Models
In this section, we introduce the generative models that we used in our experiments. These models are different both in their architectures and in the domains they were trained in. This will allow us to more thoroughly compare how their different characteristics affect the quality of the generations.
@@ II-A1 BERT family of models
Bidirectional Encoder Representations from Transformers (BERT) [5]. BERT is designed to pre-train deep bidirectional representations by conditioning on both left and right contexts in all layers. As a result, the pre-trained BERT model can be fine-tuned with just one additional output layer to create state-of-the-art models for a wide range of tasks, such as question answering [8] and natural language inference [9], without substantial task-specific architecture modifications. The model has 345M parameters.
RoBERTa Robustly Optimised BERT Pretraining Approach (RoBERTa) [10] builds upon BERT by modifying key hyper-parameters in its pre-training process, which includes training the model longer, with more data and on bigger batches. This model showed better performance than BERT across a wide range of tasks. The model has 355M parameters.
BiomedNLP-PubMedBERT BiomedNLP-BiomedBERT-large-uncased-abstract (PubMedBERT) [11] is a BERT variant fine-tuned for the biomedical domain that uses a vast corpus of biomedical texts from PubMed, making it more effective in handling the biomedical language. The model showed state-of-the-art performance on the Biomedical Language Understanding and Reasoning Benchmark (BLURB) [^1] https://microsoft.github.io/BLURB/. We want to use this model because we want to understand whether task-specific knowledge helps the model generate higher-quality text for this specific task.
@@ II-A2 T5 family of models
Text-to-Text Transfer Transformer (T5) [12] is an encoder-decoder model that frames all NLP tasks as text-to-text problems. It is pre-trained on the Colossal Clean Crawled Corpus (C4) using a denoising objective, where the model learns to reconstruct corrupted input sequences. Providing task-specific prefixes to the input text allows it to be fine-tuned for various tasks, such as translation, summarization, and question-answering. The t5-large version is made up of 770M parameters, which is a comparable size to the large models in the BERT family.
SciFive-large-Pubmed_PMC SciFive [13] is a version of the T5 model fine-tuned on scientific and biomedical literature, specifically texts taken from PubMed and PMC (PubMed Central). SciFive is excellent in tasks related to scientific text, such as document summarisation, entity extraction, and knowledge graph construction. The model is used primarily for understanding and generating scientific text. Simarly to PubMedBERT (see Section II-A1), we want to use this model to evaluate whether task-specific knowledge affects the quality of the generations.
@@ II-A3 BART
BART (Bidirectional and Auto-Regressive Transformer) [14] is a generalized pre-training model based on the Transformer architecture. The model was pre-trained using 5 techniques:
  • Token masking, in the same way as used in BERT, randomly masks tokens.
  • Sentence permutation shuffles the sentences of a document randomly.
@@
  • Token deletion takes the original sentence and randomly deletes a token from the text.
@@ II-B Data Augmentation
Data augmentation has recently been used more extensively in NLP due to more work performed in low-resource domains [7] and the increasing popularity of large neural networks that require massive amounts of training data [15]. Applying it to this field is far more challenging than in others due to the discrete nature of language, where slight modifications can significantly alter the semantics of a sentence. Notwithstanding these challenges, numerous works have applied data augmentation for NLP. These can be simple approaches, often just involving rule-based approaches that can offer minor improvements [16]. More complex approaches are often model-based and have been applied to numerous sub-fields of NLP, such as text classification [17], question answering [18], and machine translation [19].
@@ II-D Masked Language Modeling for Text Generation
The idea that BERT [23] could be used for text generation was first introduced in 2019 [6]. This approach generates some or all the tokens simultaneously by looking at a bidirectional context. Non-autoregressive text generation was primarily used in Neural Machine Translation (NMT) to speed up inference time, by decoding multiple tokens simultaneously, at the cost of sacrificed translation accuracy compared to its counterpart, autoregressive generation [24]. However, few works have explored the potential of BERT in generating text outside of machine translation [25, 26]. However these works focused on instruction following and story generations, respectively, but still focused on improving inference speed as the primary goal rather than text quality. In addition they did not use any domain-specific dataset, and did not attempt to utilize the generations in downstream tasks. In our work, we aim to expand on this by examining the conditions under which MLM models can generate better text than CLM models, thereby finding the best direction for future research in MLM for Text Generation.
@@ III-B1 Medical Dataset
III-B2 Movive Plot Synopsis Dataset
For the second task, we used the Movie Plot Synopses with Tags (MPST) Corpus [30]; the dataset comprises a corpus of 14K movie plot synopses and is associated with 70 tags per synopsis, retrieved from Wikipedia and IMDb.
@@ III-B3 Authorship Verification Dataset
The PAN Author Verification dataset [^2] https://pan.webis.de/clef23/pan23-web/author-identification.html contains texts from 112 individuals, each providing a written and spoken text of different Discourse Types on various topics and varying degrees of formality. These texts were transcribed into essays, emails, interviews, and speech transcriptions. All the texts in the dataset are from native English speakers.
@@ III-C1 MLM Models
Training We used the standard MLM training protocol to train the MLM models. This approach entails masking random tokens and then letting the model learn to predict these based on the provided context, i.e., the non-masked tokens. While typically the proportion of masked tokens in the training phase is about 15% [5], in our case, we set the proportion to 50% random tokens to let the model learn in conditions that more closely align with the generation phase.
Generation The process we implemented for generating text using the MLM models consists of two steps. First, the original text would be corrupted according to the specific corruption method described in Section III-C3 to contain masked and unmasked tokens. Second, the model would predict each [MASK] token one by one until no more mask tokens remained in the text. As shown in Figure II, the model does not need to follow a particular order to unmask each token and does so based on which prediction for each [MASK] token has the highest confidence.
@@ III-C2 CLM Models
Training To train the CLM models, we followed a similar approach as for MLM, but this time, we removed all of the masked tokens from the text and input the unmasked tokens into the model, which is then trained to generate text given the list of unmasked tokens.
Generation Generation works analogously to training: we first select the input tokens and then let the model generate the text given that context.
III-C3 Generation Approaches
To select the right tokens to be the input for the generation, we employed different approaches and assessed their impact on the performance of the generations of the models models. We used the same approaches for both MLM and CLM. The only difference was that in MLM, we replaced the tokens with [MASK], while in CLM, we removed the tokens from the input.
The following are the approaches that we used:
  • Random: The tokens were picked randomly from the input text. With this approach, we masked a percentage of tokens from 10% to 100%.  [^3] For CLM models, we retained that the evaluation of the 100% masking ratio was unfair, so we excluded that from the final evaluation. The goal of this approach was to establish the baseline for text generation and to determine the extent to which the ratio of the number of masked tokens can impact the quality of the generated text by the models.
  • Stopwords: This time, we masked all the tokens in the text that were not English stopwords. The goal of this approach was to understand whether punctuation alone could provide enough context to generate high-quality text.
@@ III-C3 Generation Approaches
  • NER: with this approach, we used SciSpacy  [^4] https://allenai.github.io/scispacy/ to detect all of the entities appearing in the text and masked the rest. This approach aimed to understand how the model can vary in its generations while maintaining the correct important entities. This generation approach was applied only to the Medical dataset, given that it was the only one with easily recognisable entities that are the core part of the text.
@@
III-D Downstream Tasks
In addition to the quality of the generations, in the first two tasks we wanted to understand their usability to train other models. In the first two tasks wanted to understand to what extent the quantitative metrics correlated with the usefulness of our texts. With the third task instead, we sought to determine which approach could generate text most resembling the original style according to a third model.
III-D1 Named Entity Recognition (NER) Task
The goal of NER is to locate and classify named entities mentioned in unstructured text into categories such as person names, organizations, locations, etc. In this task, we want to evaluate the performance of an NER model trained on the generated texts. Since the dataset is medical in this case, we will focus on the ability of the model to detect medical entities.
As Figure 2 shows, we first used SciSpacy(a NER  [^5] https://allenai.github.io/scispacy/ to detect medical entities in the synthetic text. The labels generated by SciSpacy would then be fed to the Spacy model (see SectionIII-D1) along with the synthetic text for training. Additionally, we used SciSpacy to extract the test NER labels from the test dataset. We then fed the test dataset to the trained Spacy model and compared its predictions with the NER labels extracted by SciSpacy, computing the F-1 Score.
NER model: SpaCy
@@ III-D2 Text Classification Task
Text Classification is associating labels to a corpus of texts. We perform a multilabel binary classification task. The dataset is from Authorship Verification 2023  [^6] https://pan.webis.de/clef23/pan23-web/author-identification.html. We further processed it to contain only human-generated plot synopses and reduced the number of tags to 6 to maintain only the ones appearing in most samples. This task focuses on the model’s ability to classify texts correctly. Like in the NER task, we trained a BiLSTM model on the generated texts but left the tags unchanged. The trained models were evaluated on the same testing set to measure their movie-level f-1 score, which is computed with the total number of True Positives, True Positives and False Negatives across all the categories.
With this task, we wanted to determine whether the texts generated by our models could be successfully used to train a text classification model, how much text quality mattered for this task, and which specific tokens provided enough context to generate useful text.
@@
The BiLSTM model was trained on the generated plot synopses with their corresponding tags taken from the real datasets. We then used the trained model to process the test dataset and predict the most likely tags. Finally, we compared the predictions with the true test labels and computed the movie-level F-1 score.

### v2

@@ Abstract
Large Language Models (LLMs) have revolutionised the field of Natural Language Processing (NLP) and have achieved state-of-the-art performance in almost every task in this field. However, the prevalent approach used in text generation, Causal Language Modelling (CLM), generates text sequentially from left to right and thus limits the freedom of the model by restricting when and where each token is generated. In contrast, Masked Language Modelling (MLM), primarily used for language understanding tasks, can generate tokens anywhere in the text and in any order. This paper conducts an extensive comparison of MLM and CLM approaches for text generation tasks. To do so, we pre-train several language models of comparable sizes on three different datasets, namely (1) medical discharge summaries, (2) movie plot synopses, and (3) authorship verification datasets. To assess the quality of the generations, we first employ quantitative metrics and then perform a qualitative human evaluation to analyse coherence and grammatical correctness. In addition, we evaluate the usefulness of the generated texts by utilizing them in three downstream tasks: (1) Entity Recognition, (2) Text Classification, and (3) Authorship Verification. The results show that MLM consistently outperforms CLM in text generation across all datasets, with higher quantitative scores and better coherence in the generated texts. The study also finds no strong correlation between the quality of the generated text and the performance of the models on the downstream tasks. Finally, we demonstrate that MLM has great potential for future research on text generation and provide direction for future studies in this area.
@@ I Introduction
The rise of Large Language Models (LLMs) has gained much attention in recent years. These models have shown remarkable ability in complex reasoning and generating human-like text. However, it became apparent since its very introduction that the transformer architecture has numerous limitations. In fact, the attention mechanism [1], a core component of transformers, struggles with capturing long-range dependencies and does not scale efficiently to capture longer sequences. This finding led researchers to explore various ways to overcome these limitations and introduce many transformer variants [2, 3, 4]. This collective effort also led to the exponential growth in size of these models.
@@ 5 The quick brown fox jumps over the lazy dog
This process is called Causal Language Modelling (CLM) because the probability of generating each next token is conditioned only on the tokens that appear before it in the sequence. This unidirectional approach inevitably restricts their capacity to leverage the full context of a text, as these models are prevented from incorporating or generating tokens that appear later in the sequence. Howbeit, this limitation is not present when using another approach called Masked Language Modelling (MLM).
@@
Despite their potential, MLM applications in text generation, particularly for tasks such as synthetic data generation [7], have been greatly overlooked, leaving their generative capabilities massively unexplored. In this paper, we aim to challenge this assumption in text generation by comparing CLM with MLM using models of comparable size. By exploring this potential, we aim to find new directions for this NLP area and understand the conditions under which MLM can shine at text generation.
Because very little research has been conducted on this topic, we decided to base our work on synthetic data generation to allow easier evaluation by comparing the synthetic dataset with the original one, as well as to use synthetic data for the training and evaluation of our models on downstream tasks.
@@
    Does using domain-specific knowledge for pre-training and fine-tuning increase the quality of generated texts with regard to the quantitative metrics?
@@
    Does higher quality in the texts translate to better performance on downstream tasks?
@@
    What type of tokens provide enough context to generate high-quality texts?
To answer these questions, we aimed to thoroughly evaluate and compare the ability of CLM and MLM models of comparable sizes to generate texts. The following are the main tasks of this investigation:
@@
    Fine-tune several Pre-trained Language Models (PLMs), including T5, BART, and BERT, and their variations.
@@
    Compare the generations of MLM and CLM models with standard quantitative metrics and qualitatively assess their generations.
@@
    Examine the usefulness of fine-tuned models’ generations on downstream tasks, namely Named Entity Recognition (NER), Text Classification and Authorship Verification.
@@
    Interpret the results of the evaluations and conclude on the optimal types of models and masking strategies for different text generation tasks across the three datasets.
@@ II Background and Related Work
II-A Backbone Models
In this section, we introduce the models that we used in our experiments. These models are different both in their architectures and the domains they were trained on. This will allow us to thoroughly compare how their different characteristics affect the quality of their generations.
@@ II-A1 BERT family of models
Bidirectional Encoder Representations from Transformers (BERT) [5]. BERT is designed to pre-train deep bidirectional representations by conditioning on both left and right contexts in all layers. As a result, the pre-trained BERT model can be fine-tuned with just one additional output layer to create state-of-the-art models for a wide range of tasks, such as question answering [8] and natural language inference [9], without substantial task-specific architecture modifications. The model has 345M parameters and counts multiple variations, such as the ones introduced below.
RoBERTa. Robustly Optimised BERT Pretraining Approach (RoBERTa) [10] builds upon BERT by modifying key hyper-parameters in its pre-training process, which includes training the model longer, with more data and on bigger batches. This model showed better performance than BERT across a wide range of tasks. The model has 355M parameters.
BiomedNLP-PubMedBERT. BiomedNLP-BiomedBERT-large-uncased-abstract (PubMedBERT) [11] is a BERT variant fine-tuned for the biomedical domain that uses a vast corpus of biomedical texts from PubMed, making it more effective in handling the biomedical language. The model showed state-of-the-art performance on the Biomedical Language Understanding and Reasoning Benchmark (BLURB). We decided to use this model in order to understand whether task-specific knowledge increases the quality of the generated texts for this specific task.
@@ II-A2 T5 family of models
Text-to-Text Transfer Transformer (T5) [12] is an encoder-decoder model that frames all NLP tasks as sequence-to-sequence problems. It is pre-trained on the Colossal Clean Crawled Corpus (C4) using a denoising objective, where the model learns to reconstruct corrupted input sequences. Providing task-specific prefixes to the input text allows it to be fine-tuned for various tasks, such as translation, summarization, and question-answering. The t5-large version is made up of 770M parameters, which is comparable to the size of large models in the BERT family. The following variation of T5 is used in our work.
SciFive-large-Pubmed_PMC. SciFive [13] is a version of the T5 model fine-tuned on scientific and biomedical literature, specifically texts taken from PubMed and PMC (PubMed Central). SciFive is excellent in tasks related to scientific text, such as document summarisation, entity extraction, and knowledge graph construction. The model is used primarily for understanding and generating scientific text. Similarly to PubMedBERT, we used this model to evaluate whether task-specific knowledge affects the generations’ quality.
@@ II-A3 BART
BART (Bidirectional and Auto-Regressive Transformer) [14] is a generalized pre-training model based on the Transformer architecture. This model counts 406M parameters and was pre-trained using five techniques:
  • Token masking randomly masks tokens (in the same way as for BERT).
  • Sentence permutation randomly shuffles the sentences of a document.
@@
  • Token deletion takes the original sentence and randomly deletes a token from the sequence.
@@ II-B Data Augmentation
Data augmentation has recently been used more extensively in NLP due to more work performed in low-resource domains [7] and the increasing popularity of large neural networks that require massive amounts of training data [15]. Applying it to this field is far more challenging than in others due to the discrete nature of language, where slight modifications can significantly alter the semantics of a sentence. Notwithstanding these challenges, numerous works have applied data augmentation for NLP. These can be simple approaches, often just involving rule-based techniques that can offer minor improvements [16]. More complex approaches are often model-based and have been applied to numerous sub-fields of NLP, such as text classification [17], question answering [18], and machine translation [19].
@@ II-D Masked Language Modeling for Text Generation
The idea that BERT [23] could be used for text generation was first introduced in 2019 [6]. This approach generates some (or all) the tokens simultaneously by looking at a bidirectional context. Non-autoregressive text generation was primarily used in Neural Machine Translation (NMT) to speed up inference time, by decoding multiple tokens simultaneously, at the cost of sacrificed translation accuracy compared to its counterpart, autoregressive generation [24]. However, very few works have explored the potential of BERT in generating text outside of machine translation [25, 26]. In fact, these works on instruction following and story generations still focused on improving inference speed rather than generation quality. In addition, they did not use any domain-specific dataset and did not attempt to utilize the generations in downstream tasks. In our work, we aim to expand on this by examining the conditions under which MLM models can generate better texts than CLM models, thereby finding the best direction for future research in MLM for text generation.
@@ III-B1 Medical Dataset
III-B2 Movie Plot Synopsis Dataset
For the second task, we used the Movie Plot Synopses with Tags (MPST) Corpus [30]. This dataset comprises a corpus of 14K movie plot synopses and is associated with 70 tags per synopsis, all retrieved from Wikipedia and IMDb.
@@ III-B3 Authorship Verification Dataset
The PAN Author Verification dataset [^1] https://pan.webis.de/clef23/pan23-web/author-identification.html contains texts from 112 individuals, each providing a written and spoken text of different Discourse Types on various topics and varying degrees of formality. These texts were transcribed into essays, emails, interviews, and speech transcriptions. All the texts in the dataset were obtained from native English speakers.
@@ III-C1 MLM Models
Training. We used the standard MLM training protocol to train the MLM models. This approach entails masking random tokens and then letting the model learn to predict these based on the provided context, i.e., the non-masked tokens. While typically the proportion of masked tokens in the training phase is about 15% [5], in our case, we set the proportion to 50% random tokens to let the model learn in conditions that more closely align with the generation phase. An example is displayed in Figure 1.
Generation. The process we implemented for generating texts using the MLM models consists of two steps. First, the original text would be corrupted according to the specific corruption method described in Section III-C3 to contain masked and unmasked tokens. Second, the model would predict each [MASK] token one by one until no more mask tokens remained in the text. As shown in Table II, the model does not need to follow a particular order to unmask each token. Instead, it predicts for each [MASK] token in the order of highest confidence.
Figure 1: Examples of Inputs and Outputs for MLM models.
@@ III-C2 CLM Models
Training. In order to train the CLM models, we followed a similar approach to that of MLMs, but removed all of the [MASK] tokens from the corrupted text and only input the unmasked tokens into the model. The latter is then trained to generate text given the list of unmasked tokens. An example can be found in Figure 2.
Generation. The generation works analogously to training: we first select the input tokens and then let the model generate the text given that context.
Figure 2: Examples of Inputs and Outputs for CLM models.
III-C3 Masking Techniques
To select the best tokens to be the input for the generations, we employed different approaches and assessed their impact on the quality of the generations. We used the same approaches for both MLM and CLM models with the only exception being that [MASK] tokens were removed for CLM models as explained above.
The masking approaches employed are listed below:
  • Random: The tokens were picked randomly from the input text. With this approach, we masked a percentage of tokens from 10% to 100%  [^2] For CLM models, we retained that the evaluation of the 100% masking ratio was unfair, so we excluded that from the final evaluation.. The aim was to establish the baseline for text generation and to determine the extent to which the ratio of the number of masked tokens can impact the quality of the generated text.
  • Stopwords: All tokens that are not English stopwords were masked. The aim was to understand whether punctuation alone could provide enough context to generate high-quality text.
@@ III-C3 Masking Techniques
  • NER: In this approach, we used SciSpacy [^3] https://allenai.github.io/scispacy/ to detect all entities in the text and mask the rest. This approach aimed to understand how the model can vary in its generations while maintaining the correct important entities. This generation approach was applied only to the Medical dataset, given that it was the only one with meaningful and easily recognisable entities.
III-D Downstream Tasks
In the first two tasks, we aimed to assess the quality of the generated text and its usability for training other models. We wanted to determine how quantitative metrics related to the usefulness of our texts. In the third task, we aimed to identify the approach that could produce the most similar texts to the original style according to a pre-trained model.
III-D1 Named Entity Recognition (NER) Task
By solving the NER task, we aim to locate and classify named entities mentioned in an unstructured text into categories such as person names, organizations, locations, etc.
In this task, we evaluate the performance of a NER model trained on the generated texts to detect medical entities.
@@ III-D1 Named Entity Recognition (NER) Task
As shown in Figure 3, we first used a pre-trained NER model, named SciSpacy, to detect medical entities in the synthetic texts. The labels generated by SciSpacy would then be fed to the Spacy model along with the synthetic texts for training. Additionally, we used SciSpacy to extract entity labels from the test dataset. We then fed the testing data to the trained Spacy model and compared its predictions with the labels extracted by SciSpacy, to compute the F1 score.
@@ III-D2 Text Classification Task
Text classification aims to associate labels to each text of a given corpus. In our case, we perform multi-label classification, meaning that each text may have multiple correct labels. The dataset is from the MPST Corpus, which we further processed to contain only human-generated plot synopses and reduced the set of tags to the six most frequent ones. This task focuses on the model’s ability to classify texts correctly. As with Named Entity Recognition, we trained a BiLSTM model on the generated texts while keeping the tags unchanged. The trained models were evaluated on the same testing set to measure their movie-level F1 score, which we computed from the total number of True Positives, False Positives and False Negatives across all categories.
@@
By following the downstream pipeline displayed in Figure 4, we aim to determine whether the texts generated by our models could be successfully used to train a text classification model. More specifically, we intent to demonstrate how much text quality matters for this task and which types of tokens provide enough context to generate useful sequences.
The BiLSTM model was trained on the generated plot synopses with their corresponding tags taken from the true dataset. We then used the trained model to process the test dataset and predict the most likely tags. Finally, we compared the predictions with the true testing labels and computed the movie-level F1 score.